how would i set this problem up? a chemist has a 25% hydrochloric acid solution and a 50% hydrochloric acid solution. how many liters of each should she mix to get 10 liters of a hydrochloric acid solution with 40% acid concentration?
Answer by richwmiller(17219)

a+b=10,
0.25*a+0.5*b=0.4*10
a=10-b
0.25*(10-b)+0.5*b=4
2.5-0.25b+0.5*b=4
0.25*b=1.5
b=6
a=10-b
a=4 b=6
check
0.25*4+0.5*6=0.4*10
1+3=4
4=4
ok